Problem 1: Solve the initial value problems: a 2y – 3y +y=0 y(0) = 2, 7(0) = 1 by + y - 6y = 0 y(0) = -1, y(0) = 2 cy +
Answer #1

@son: Given Intial value Problem 24-Bylty=0y60)=2, y(o)=% -O The Auniliary equation of yl-3yltyao is fimi zo =) a m-3m +12Given that y+y-by=0, 910)=-1, y(0)=2- The Aumiliary equation of ytyl-by zo is from=0=) i tm-t=0 =) m +3m- am-6=0 = mint 3